# Nontrivial twisted bilayer chiral excitonic systems: Förster coupling and related Hall effect

### v1

@@ Abstract
In bilayers composed of monolayers of two-dimensional semiconductors with arbitrary twisting angles, an effective excitonic system can be established based on the valley excitons of each layer. Instead of direct interlayer hopping, a pronounced Förster coupling connects intralayer excitons in different layers, which modifies the energy of the monolayer excitonic system. Surprisingly, two new types of Hall effect recently discovered in electric systems-the time-reversal even charge Hall effect (TREHE) and the crossed nonlinear dynamical Hall effect (CNDHE)-can occur in these systems due to the chiral structure induced by the twist of two layers. We illustrate this point by employing a general theoretical framework to establish the effective excitonic Hamiltonians in twisted homobilayer systems composed of transition metal dichalcogenides (TMD) and black phosphorus (BP), calculating the Hall effect in these systems as examples. Regarding the TREHE, our calculations show that it depends on the anisotropy of excitons, resulting in a negligible Hall conductivity for TMD without strain due to the isotropic excitons, but a large Hall conductivity for TMD with chiral symmetric excitons induced by strain. Meanwhile, BP exhibits a finite Hall conductivity due to its high anisotropy. As for the CNDHE, only the specific value of twisted angles can induce a significant Hall conductivity in BP and TMD with strain. This work uncovers nontrivial physical properties of twisted bilayer chiral excitonic systems and provides new opportunities for valley exciton optoelectronics in the layer degree of freedom.
@@
The study of emerging layered two-dimensional (2D) semiconductors has increasingly become one of the most crucial fields in condensed matter physics [1, 2, 3, 4, 5, 6]. Several of these materials can serve as a reliable platform to explore exciton physics due to strong light-matter interactions [2, 3, 4]. In transition metal dichalcogenides (TMDs), tightly bound Wannier excitons can emerge around the degenerate ±$\pm K$ valleys at the corners of the Brillouin zone (BZ), with optical selection rules for manipulating the valley pseudospin [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The small Bohr radius $\sim O(1)$ nm [17, 18, 19] indicates a significant electron-hole (e-h) Coulomb exchange at finite center-of-mass (COM) momentums [13, 20, 21, 22]. This coupling between the exciton’s valley pseudospin and COM degrees of freedom results in the splitting of the exciton dispersion into two branches having linearly polarized optical dipoles, longitudinal ($L$) and transverse ($T$) to exciton momentum, respectively [13, 21, 23, 24, 25, 26]. Recently, researchers have studied the exotic properties of ground-state excitons in monolayer TMDs on patterned substrates [27, 28] due to the massless $L$ branch’s sensitive dependence on the surrounding dielectric [29, 30]. Also, the $C_{3}$ symmetry of monolayer TMDs can be broken by applying the strain into it, which leads to a non-zero contribution of short-range e-h exchange between excitons in different valleys, causing linearly dispersing Dirac saddle points in the light cone [13, 21].
@@
Otherwise, in highly anisotropic materials such as the black phosphorus (BP), which is a potentially ideal platform for high-performance devices like a field-effect transistor [31] since monolayer BP is a semiconductor with a predicted direct bandgap of $1\sim 2\,\mathrm{eV}$ at the $\Gamma$ point in the BZ [32, 33, 34]. The highly anisotropic band structure of the lowest conduction band and highest valence band [32, 33, 34] leads to anisotropic valley excitons [33, 35] corresponding with almost linearly polarized optical dipole instead of circularly polarized one for ±$\pm K$ isotropic valley excitons in monolayer TMDs [21, 22, 23].
@@
In contrast to the extensively studied electrical behaviors in twisted bilayer and multi-layer systems [2, 5, 6, 7], or excitonic systems that rely on the interaction between interlayer and intralayer excitons [3, 4, 21, 37], the effective system related to Förster coupling [38], specifically the Coulomb exchange between electron-hole (e-h) pairs, which can transfer an exciton or e-h pair non-locally between intralayer excitons in different layers, has seldom been considered. This interaction enables out-of-plane propagation of excitons in a stack of layers, and its long-range nature allows it to work even when charge hopping is completely quenched by spacer layers or crystalline misalignment. Also, previous studies have proved that cross-dimensional valley excitons can arise from this coupling in arbitrarily twisted stacks of monolayer semiconductors [39].
Just as the anomalous Hall effect for electrons that results from the Berry curvature in momentum space [40], the exciton Hall effect can also arise due to the exciton Berry curvature [41, 42]. Recently, two new types of Hall effect have been discovered in twisted bilayer chiral electric systems, i.e., the time-reversal even charge Hall effect (TREHE) based on the spectral chiral symmetry [43] and the crossed nonlinear dynamical Hall effect (CNDHE) in twisted bilayers, induced by out-of-plane alternating-current (AC) electric fields [44, 45]. These findings uncover novel Hall physics in chiral structures and pave the way for layertronics, a promising avenue for exploring the quantum nature of the layer degree of freedom and discovering exciting effects. The question that naturally arises is whether these effects can also occur in twisted bilayer chiral excitonic systems and be experimentally observed.
In this paper, we investigate the properties of twisted bilayer chiral excitonic systems that arise from intralayer excitons in each layer and the Förster coupling connecting them across different layers. Using twisted homobilayer TMDs and BP as examples, we build a general method to derive effective excitonic Hamiltonians from electron-hole Coulomb exchange interactions and demonstrate that two newly discovered Hall effects, i.e., the TREHE and the CNDHE, in electronic systems also exist in excitonic systems. Our calculations reveal that the twist-induced chiral structure, material anisotropy, and twist angle significantly influence these effects. These findings reveal nontrivial properties and present new opportunities for valley exciton optoelectronics.
@@
Figure 1: (color online) (a) Left panel: Schematic of an arbitrarily twisted homobilayer BPs, where hBN spacers quench the charge hopping. Right panel: Illustration of electron-hole Coulomb exchange between valley excitons in the same or different layers, showing approximate band edges near high-symmetry points. (b) Similar plot for twisted homobilayer TMDs with strain (applied along $x$ direction with strain extent $\epsilon_{st}\sim 1\%$ in the bottom layer). (c) Left panel: Excitonic dispersions for the effective excitonic Hamiltonian (26) of BP with different twisted angles $\theta$. Right panel: Anisotropic polarization distribution of the optical dipole for two bands in momentum space. (d) Similar plots for the effective excitonic Hamiltonian (17) of TMD with and without strain in untwisted situation. For clarity, only the polarization distribution $p_{n}$ of the optical dipole for the marked band $\varepsilon_{n}$ are shown.
I Effective excitonic Hamiltonians of twisted homobilayer TMDs and BP
@@
As previously stated, the effective excitonic Hamiltonian in twisted bilayer structures is primarily based on intralayer valley excitons in the two layers. In the case of monolayer situations under the low-energy approximation, there are typically at least two non-equivalent valleys connected by rotational or time-reversal symmetry in the first BZ, such as the ±$\pm K$ valley in monolayer TMDs [2, 7] (Fig. 1(b)), resulting in a $4\times 4$ effective Hamiltonian for intralayer excitons, as discussed below. However, when only a single valley is present, such as the $\Gamma$ point in monolayer BP (Fig. 1(a)), the Hamiltonian reduces to a simple $2\times 2$ matrix. In the following, we present a method for obtaining the effective excitonic Hamiltonian through a general calculation of the e-h Coulomb exchange interaction, and use it to acquire the excitonic Hamiltonian for these two materials.
I.1 General description the effective excitonic Hamiltonian
@@
The construction of this Hamiltonian relies on two types of interaction: the e-h exchange interaction between intralayer excitons of the same or different valleys in the same layer and the Förster coupling that connects excitons in different layers, as illustrated in Fig. 1. For a specific layer $l$ ($l=t,b$ in this paper, $t/b$ means top/bottom layer, as shown in Fig. 1(a)), the basis of valley excitons can be expressed as $\left(\left|l,\boldsymbol{k}\right\rangle_{\alpha},\left|l,\boldsymbol{k}\right\rangle_{\beta}\right)$, $\alpha,\beta$ means two different valley indexes with the in-plane COM momentum $\boldsymbol{k}=\left(k\cos\varphi,k\sin\varphi\right)$, and generally the effective excitonic Hamiltonian in this situation can be expressed as
@@
$J_{\lambda,\lambda^{\prime}}^{l}$$\approx$ $\psi_{\lambda}^{l}\left(0\right)\left[\psi_{\lambda^{\prime}}^{l}\left(0\right)\right]^{\ast}\frac{V\left(\boldsymbol{k}\right)}{4}\times$
@@
I.2 Excitonic Hamiltonians of TMDs
For the TMD cases, there are two inequivalent valleys as $\alpha\equiv\boldsymbol{K}$ and $\beta\equiv-\boldsymbol{K}$, connected by the time-reversal symmetry. It has been shown that the excitonic dispersion is isotropic for the first order of $\boldsymbol{k}$ in monolayer TMDs [13, 21, 23], these two properties lead to $H_{k}^{l}=\frac{\hbar^{2}k^{2}}{2m_{ex}}=H_{k}$ and $a_{B,\lambda}^{x,l}=a_{B,\lambda}^{y,l}=a_{B,\lambda^{\prime}}^{x,l^{\prime}}=a_{B,\lambda^{\prime}}^{y,l^{\prime}}=a_{B}$. Without loss of generality, we let $\theta_{b}=0$ and $\theta_{t}=\theta$ and replace the twist angle $\theta$ in momentum space into the real space one, i.e., $\theta\rightarrow-\theta$ in the rest of the paper. The effective Hamiltonian $H_{\mathrm{TMD}}$ of valley excitons in the twisted homobilayer TMD can be written as
@@ I.2 Excitonic Hamiltonians of TMDs
which is independent of the twisted angle $\theta$. $m_{ex}\approx m_{e}$ is the effective exciton mass in monolayer TMDs. Here $K=4\pi/3a$, $a$ being TMD’s lattice constant, and $J\sim 1\,\mathrm{eV}$ can be extracted from first principle wavefunctions and exciton spectrum [7, 21, 23]. The isotropic properties of the excitonic bands in momentum space are evident from Fig. 1(c).
If there is a strain in the bottom layer of twisted bilayer TMDs, it can be transferred into the top layer with some loss [46, 47, 48]. This mechanism leads to chiral symmetric excitons in twisted bilayer TMDs, where the short-range e-h exchange contributes to the coupling between excitons in different valleys of the same layer. This results in the non-diagonal term of $H_{\mathrm{intra}}^{l}$ becoming $J_{0}^{l}-J\frac{k}{K}e^{\pm i\left(\theta_{l}+2\varphi\right)}$ [36]. In our consideration, $J_{0}^{b}\approx-6\,\mathrm{meV}$ arises from the strain along $x$ direction with an extent of $\epsilon_{st}=1\%$ [21], and $J_{0}^{t}\approx J_{0}^{b}\frac{\cos 2\theta+\sin 2\theta}{2}$ is twisted-angle dependent [36]. The related excitonic dispersions, with and without strain in untwisted situation, are displayed in Fig. 1(d), which also shows the polarization of the optical dipole for the band $\varepsilon_{n=1,3}$. The anisotropic polarization distribution of optical dipole in momentum space indicates that the chiral symmetry in real space preserves in the effective excitonic Hamiltonian when strain is applied.
I.3 Excitonic Hamiltonian of BP
The excitonic Hamiltonian of BP can be obtained in the basis of $\{\left|b,\boldsymbol{k}\right\rangle_{\Gamma},\left|t,\boldsymbol{k}\right\rangle_{\Gamma}\}$ by using the same approach as for TMDs. Since there is only one valley ($\Gamma$ point) per layer under the low-energy approximation, i.e., $\alpha=\beta\equiv\Gamma$, this Hamiltonian takes a simple $2\times 2$ form as [36]
@@ I.3 Excitonic Hamiltonian of BP
The excitonic dispersions of $H_{\mathrm{BP}}$ has the from
@@
Unlike in TMDs, $\theta$ affects the excitonic properties of $H_{\mathrm{BP}}$, which can be seen from the highly anisotropic dispersion at $\theta=0$ and $\pi/6$, also the highly anisotropic polarization distribution of optical dipole, as shown in Fig. 1(c).
Figure 2: (color online) (a) Schematic illustration of the time reversal even Hall effect for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ (Eq. 28) for $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) Similar plot for $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 2\,\mathrm{eV}$. $\tau=1\,\mathrm{ps}$ for all plots [43].
II The time-reversal even and crossed nonlinear dynamical Hall effect in those systems
II.1 The time-reversal even Hall effect (TREHE)
The non-zero TREHE requires chiral symmetry in real space [43], which is naturally satisfied by the twisted bilayer structure, as shown in Fig. 2(a). In twisted bilayer chiral excitonic systems, this effect can be defined as similar as the one for twisted bilayer electric systems [43], which comes from the semi-classical calculation for the current density of each layer

### v2

# Chiral excitonic systems in twisted bilayers from Förster coupling and unconventional excitonic Hall effects

@@ Abstract
In twisted bilayer semiconductors with arbitrary twisting angles, a chiral excitonic system can arise from the interlayer electron-hole Coulomb exchange interaction (Förster coupling) that hybridizes the anisotropic intralayer excitons from individual layers. We present a general framework for the effective exciton Hamiltonian taking into account the electron-hole Coulomb exchange, using twisted homobilayer systems composed of transition metal dichalcogenides or black phosphorus as examples. We demonstrate that such chiral excitonic systems can feature unconventional Hall (Nernst) effects arising from quantum geometric properties characteristic of the layer hybridized wavefunctions under the chiral symmetry, for example, the time-reversal even layer Hall counter flow and the crossed nonlinear dynamical Hall effect, when mechanical and statistical force (temperature or density gradient) drives the exciton flow.
@@
Layered two-dimensional (2D) semiconductors have provided a platform for the various frontiers of condensed matter physics [1, 2, 3, 4, 5, 6]. In particular, they offer exciting opportunities to explore exciton physics and strong light-matter interactions [2, 3, 4]. In transition metal dichalcogenides (TMDs), tightly bound Wannier excitons are formed around the degenerate ±$\pm K$ valleys at the corners of the Brillouin zone (BZ), where optical selection rules make possible manipulation of the valley pseudospin [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The small Bohr radius $\sim O(1)$ nm [17, 18, 19] underlies a significant electron-hole (e-h) Coulomb exchange at finite center-of-mass (COM) momentum [13, 20, 21, 22]. This effective coupling between the exciton’s valley pseudospin and COM degrees of freedom results in the splitting of the exciton dispersion into two branches having linearly polarized optical dipoles, longitudinal ($L$) and transverse ($T$) to exciton momentum, respectively [13, 21, 23, 24, 25, 26]. Recent studies have found exotic properties of excitons in monolayer TMDs on patterned substrates [27, 28] due to the massless $L$ branch’s sensitive dependence on the surrounding dielectric [29, 30]. Additionally, applying strain to monolayer TMDs can break the rotational symmetry. This symmetry breaking results in a non-zero contribution from short-range e-h exchange interactions between excitons in different valleys, leading to anisotropy and the formation of linearly dispersing Dirac saddle points within the light cone [13, 21].
@@
Black phosphorus (BP) is an example of highly anisotropic 2D semiconductors. The monolayer has a direct bandgap of $1\sim 2\,\mathrm{eV}$ at the $\Gamma$ point in the BZ [31, 32, 33], suitable for hosting field-effect transistors as well [34]. The highly anisotropic band structure of the band edges leads to anisotropic intralayer excitons [32, 35], which exhibit linearly polarized optical dipoles, in contrast to the circularly polarized one for isotropic valley excitons in monolayer TMDs [21, 22, 23].
@@
In layered structures, exciton’s layer degree of freedom can also be explored for novel functionalities, which is typically introduced by interlayer carrier tunneling that leads to the hybridization of intra- and inter-layer exciton species [3, 4, 21, 37]. On the other hand, Förster coupling [38], i.e., the e-h Coulomb exchange, provides another channel to introduce nontrivial layer structures for excitons. This interaction can non-locally transfer an intralayer exciton between different layers [39, 40], enabling exciton propagation in the out-of-plane direction. Its long-range nature allows it to function even when carrier interlayer tunneling is quenched by spacer layers. Previous studies have shown that cross-dimensional valley excitons can arise from this coupling in arbitrarily twisted stacks of monolayer semiconductors [41, 42].
Like electrons where the band geometric quantities have been extensively explored in transport phenomena [43], excitons can also have such properties from the dependence of its internal degree of freedom on the COM momentum, giving rise to novel exciton transport driven by the external mechanical and statistical force (temperature or density gradient) [44, 45]. In layered structures, this possibility can be rooted in the layer degree of freedom in the exciton wavefunction which can be engineered by twisting in the layered structure. Recently, two new types of Hall effect have been discovered in chiral electronic systems in twisted bilayers, i.e., the time-reversal even layer Hall counter flow (TREHCF) [46] and the crossed nonlinear dynamical Hall effect (CNDHE) [47, 48]. These findings uncover novel Hall physics in chiral electronic systems that can be straightforwardly engineered by twisting. The question that naturally arises is whether chiral excitonic systems can be similarly engineered.
In this paper, we show that the structural chirality from twisting in a layered structure can lead to a chiral exciton system when Förster coupling hybridizes intralayer excitons from the individual layers. In the limit where carrier interlayer tunneling is quenched, we present a general formulation of the effective exciton Hamiltonian taking into account the e-h Coulomb exchange interactions, using twisted homobilayer TMDs and BP as examples. The excitonic chirality requires the presence of anisotropy that breaks the continuous in-plane rotational symmetry, which is satisfied by the strong anisotropic exciton in BP. In the TMDs case, the required anisotropy can be provided by either trigonal warping or strain, and we considered the latter as an example. We demonstrate that such chiral excitonic systems can also feature unconventional Hall (Nernst) effects arising from quantum geometric properties characteristic of the layer hybridized wavefunctions under the chiral symmetry, for example, the TREHCF and the CNDHE, when mechanical and statistical force (temperature or density gradient) drives the longitudinal exciton flow. These findings point to exciton-based optoelectronic functionalities exploiting the layer quantum degree of freedom in twisted structures.
@@
Figure 1: (color online) (a) Left panel: Schematic of an arbitrarily twisted homobilayer TMDs with strain (applied along $x$ direction with strain extent $\epsilon_{st}\sim 1\%$ in the bottom layer), where hBN spacers quench the charge hopping. Right panel: Illustration of electron-hole Coulomb exchange between valley excitons in the same or different layers, showing approximate band edges near high-symmetry points. (b) Similar plot for twisted homobilayer BP. (c) Left panel: Excitonic dispersions for the effective exciton Hamiltonian (17) of TMDs with and without strain in untwisted situation. Right panel: Polarization distribution of the optical dipole for two bands in momentum space. For clarity, only the polarization distribution $p_{n}$ of the optical dipole for the marked band $\varepsilon_{n}$ are shown. (d) Similar plots for the effective exciton Hamiltonian (26) of BP with different twisted angles $\theta$.
I Effective exciton Hamiltonians of twisted homobilayer TMDs and BP
@@
As previously stated, the effective exciton Hamiltonian in twisted bilayer structures is primarily based on intralayer valley excitons in the two layers. In the case of monolayer situations under the low-energy approximation, there are typically at least two non-equivalent valleys connected by rotational or time-reversal symmetry in the first BZ, such as the ±$\pm K$ valley in monolayer TMDs [2, 7] (Fig. 1(a)), resulting in a $4\times 4$ effective Hamiltonian for intralayer excitons, as discussed below. However, when a single high symmetry point is present, such as the $\Gamma$ point in monolayer BP (Fig. 1(b)), the Hamiltonian reduces to a simple $2\times 2$ matrix. In the following, we present a method for obtaining the effective exciton Hamiltonian through a general calculation of the e-h Coulomb exchange interaction, and use it to acquire the exciton Hamiltonian for these two materials.
I.1 General description the effective exciton Hamiltonian
@@
The construction of this Hamiltonian relies on two types of interaction: the e-h exchange interaction between intralayer excitons of the same or different valleys in the same layer and the Förster coupling that connects intralayer excitons in different layers, as illustrated in Fig. 1. For a specific layer $l$ ($l=t,b$ in this paper, $t/b$ means top/bottom layer, as shown in Fig. 1(a)), the basis of valley excitons can be expressed as $\left(\left|l,\boldsymbol{k}\right\rangle_{\alpha},\left|l,\boldsymbol{k}\right\rangle_{\beta}\right)$, $\alpha,\beta$ means two different valley indexes with the in-plane COM momentum $\boldsymbol{k}=\left(k\cos\varphi,k\sin\varphi\right)$, and generally the effective exciton Hamiltonian in this situation can be expressed as
@@
$J_{\lambda,\lambda^{\prime}}^{l}$$\approx$ $\left[\psi_{\lambda}^{l}\left(0\right)\right]^{\ast}\psi_{\lambda^{\prime}}^{l}\left(0\right)\frac{V\left(\boldsymbol{k}\right)}{4}\times$
@@
I.2 The exciton Hamiltonian of TMDs
For the TMD cases, there are two inequivalent valleys as $\alpha\equiv\boldsymbol{K}$ and $\beta\equiv-\boldsymbol{K}$, connected by the time-reversal symmetry. It has been shown that the exciton dispersion is isotropic for the first order of $\boldsymbol{k}$ in monolayer TMDs [13, 21, 23], these two properties lead to $H_{k}^{l}=\frac{\hbar^{2}k^{2}}{2m_{ex}}=H_{k}$ and $a_{B,\lambda}^{x,l}=a_{B,\lambda}^{y,l}=a_{B,\lambda^{\prime}}^{x,l^{\prime}}=a_{B,\lambda^{\prime}}^{y,l^{\prime}}=a_{B}$. Without loss of generality, we let $\theta_{b}=0$ and $\theta_{t}=\theta$ and replace the twist angle $\theta$ in momentum space into the real space one, i.e., $\theta\rightarrow-\theta$ in the rest of the paper. The effective Hamiltonian $H_{\mathrm{TMD}}$ of valley excitons in the twisted homobilayer TMD can be written as
@@ I.2 The exciton Hamiltonian of TMDs
which is independent of the twisted angle $\theta$. $m_{ex}\approx m_{e}$ is the effective exciton mass in monolayer TMDs. Here $K=4\pi/3a$, $a$ being TMD’s lattice constant, and $J\sim 1\,\mathrm{eV}$ can be extracted from first principle wavefunctions and exciton spectrum [7, 21, 23]. The isotropic properties of the exciton bands in momentum space are evident from Fig. 1(c).
If there is a strain in the bottom layer of twisted bilayer TMDs, it can be transferred into the top layer with some loss [49, 50, 51]. This mechanism leads to chiral symmetric excitons in twisted bilayer TMDs, where the short-range e-h exchange contributes to the coupling between excitons in different valleys of the same layer. This results in the non-diagonal term of $H_{\mathrm{intra}}^{l}$ becoming $J_{0}^{l}-J\frac{k}{K}e^{\pm i\left(\theta_{l}+2\varphi\right)}$ [36]. In our consideration, $J_{0}^{b}\approx-6\,\mathrm{meV}$ arises from the strain along $x$ direction with an extent of $\epsilon_{st}=1\%$ [21], and $J_{0}^{t}\approx J_{0}^{b}\frac{\cos 2\theta+\sin 2\theta}{2}$ is twisted-angle dependent [36]. The related exciton dispersions, with and without strain in untwisted situation, are displayed in Fig. 1(c), which also shows the polarization of the optical dipole for the band $\varepsilon_{n=1,3}$. The anisotropic polarization distribution of optical dipole in momentum space indicates that the excitonic chirality appears in twisted bilayer TMDs when strain is applied.
I.3 The exciton Hamiltonian of BP
The exciton Hamiltonian of BP can be obtained in the basis of $\{\left|b,\boldsymbol{k}\right\rangle_{\Gamma},\left|t,\boldsymbol{k}\right\rangle_{\Gamma}\}$ by using the same approach as for TMDs. Since the band edge only appear in the $\Gamma$ point per layer under the low-energy approximation, i.e., $\alpha=\beta\equiv\Gamma$, this Hamiltonian takes a simple $2\times 2$ form as [36]
@@ I.3 The exciton Hamiltonian of BP
The exciton dispersions of $H_{\mathrm{BP}}$ has the from
@@
Unlike in TMDs, $\theta$ affects the excitonic properties of $H_{\mathrm{BP}}$, which can be seen from the highly anisotropic dispersion at $\theta=0$ and $\pi/6$, also the highly anisotropic polarization distribution of optical dipole, as shown in Fig. 1(d).
Figure 2: (color online) (a) Schematic illustration of the time reversal even layer Hall (Nernst) counter flow for excitons in twisted bilayer systems. (b) Numerical calculation of $\sigma_{H}^{\mathrm{sys}}$ and $\alpha_{H}^{\mathrm{sys}}$ (Eq. 28) for $H_{\mathrm{TMD}}$ (Eq. 17) when strain is applied, with respect to the twisted angle $\theta$ and the temperature $T$, respectively. $\mu\approx 1.69\,\mathrm{eV}$. (c) Similar plot for $H_{\mathrm{BP}}$ (Eq. 26), with $\mu\approx 2\,\mathrm{eV}$. $\tau=1\,\mathrm{ps}$ for all plots [46].
II The time-reversal even layer Hall (Nernst) counter flow and crossed nonlinear dynamical Hall (Nernst) effect
II.1 The time-reversal even layer Hall (Nernst) counter flow (TREHCF)
The non-zero TREHCF requires chiral symmetry in real space [46], which is naturally satisfied by the twisted bilayer structure, as shown in Fig. 2(a). In twisted bilayer chiral excitonic systems, this effect can be defined as similar as the one for twisted bilayer electric systems [46], which comes from the semi-classical calculation for the current density of each layer